OpenStudy (anonymous):

I need to find the derivative to 8/sqrt(4+3x)

OpenStudy (walters):

\[=8\left( 4+3x \right)^{\frac{ -1 }{ 2 }}\] =\[\frac{ -1 }{ 2 }(8)(4+3x)^{\frac{ -3 }{ 2}}(3)=\frac{ -12 }{ \sqrt[3]{4+3x} }\]